OpenStudy (anonymous):

Indicate the equation of the line that is the perpendicular bisector of the segment with endpoints (4, 1) and (2, -5). Place the symbols on the grid to complete the equation.

OpenStudy (anonymous):

perpendicular bisector passes through mid point. So, the mid point = (3,-2) since, lines are perpendicular. product of slopes is -1 Slope of given line = y2-y1 / x2-x1 = 2-4 / -5-1 = -2/-6 = 1/3 So, the slope of bisector = -3 Equation of bisector can be found out using equation of a line having slope m and passing through (x1,y1) : y-y1 = m (x-x1) y-(-2) = -3(x-3) y+2 = -3x + 9 3x+y-7 = 0

OpenStudy (anonymous):

Thank you!